An open-end mutual fund has the following stocks:
Stock Shares Stock Price
A 6,000 $98
B 33,000 19
C 4,600 89
D 82,500 12

Assume there are 50,000 shares of the mutual fund, and liabilities of $110,000. Assume the fund is sold with a 6.25 percent front-end load. What is the offering price of the fund? (Do not round intermediate calculations. Round your answer to 2 decimal places. Omit the "$" sign in your response.)

  Offering price $   
Answer #1

Total assets = (6,000 × $98) + (33,000 × $19) + (4,600 × $89) + (82,500 × $12) = $2,614,400

NAV = ($2,614,400 - $110,000) / 50,000 = $50.09

Offering price = $50.09 / (1 – .0625) = $52.43